Find the volume of the largest cube that can be cut from a whose a log of cross section whose radius is 12 centimeters
1
Expert's answer
2021-10-25T16:05:43-0400

Consider a square inscribed in a circle. The radius of a wood log or a circle is equal to "1\/2" of its diagonal "d." By Pythagorean Theorem, find the side of the square "a"


"a^2+a^2=d^2"

"a=d/\sqrt{2}"

Substitute


"a=\dfrac{2(12\ cm)}{\sqrt{2}}=12\sqrt{2}\ cm"

If the length of the log is no less than "12\sqrt{2}\ cm," then the volume of the largest cube that can be cut from a whose a log is


"V=a^3=(12\sqrt{2}\ cm)^3=3456\sqrt{2}\ cm^3"